OpenStudy (anonymous):

to solve for X i could use this = (31) + 6 or x = -(31) + 6 problem:x^2 - 12x + 5 = 0 you may just say yes or no thx

OpenStudy (anonymous):

close

OpenStudy (anonymous):

want a hint?

OpenStudy (anonymous):

well yes

OpenStudy (anonymous):

you forgot to take the square root of 31

OpenStudy (anonymous):

once you had \((x-6)^2=31\) the next step was \(x-6=\pm\sqrt{31}\) not \(x-6=\pm31\)

OpenStudy (anonymous):

o so il use 2 instead of 6

OpenStudy (anonymous):

no no you had it correct, you just forgot the square root

OpenStudy (anonymous):

it is not \(-31+6\) but rather \(-\sqrt{31}+6\) and \(\sqrt{31}+6\)

OpenStudy (anonymous):

you don't know the square root of 31, so just leave it in your answer